OpenStudy (anonymous):

Find the missing measure for a right circular cone given the following information. Find h if r = 5 and V = 100pi 4 12 30

OpenStudy (anonymous):

Thanks! (:

OpenStudy (anonymous):

I here for u(:

OpenStudy (caozeyuan):

V=(1/3)pi*r^2*h so 300=r^2*h h=300/25, so it's 12